# ManipDreamer3D : Synthesizing Plausible Robotic Manipulation Video with Occupancy-aware 3D Trajectory

### v1

@@ Introduction
To evaluate the quality of generated videos, we use common video qualities including FVD (Unterthiner et al. 2018), SSIM (Wang et al. 2004), and PSNR (Hore and Ziou 2010), as well as VBench (Huang et al. 2024b) metrics. Aside from that, we also evaluate our model with trajectory error, which reveals how well the model follows the condition. Following  (Fu et al. 2025), we evaluated this metric for the robot end-effector and the target object, respectively.
@@
The quantitative comparisons of video generation quality are presented in Tab. 2 and Tab. 3. Our ManipDreamer3D demonstrates superior performance across both evaluation frameworks.
In common video metrics (Tab. 2), our SVD variant achieves best-in-class scores with FVD (143.33), PSNR (22.75), and SSIM (0.807), while significantly reducing trajectory errors (17.40 for robot, 18.77 for object) compared to SVD-based competitors including DragAnything (Wu et al. 2024) and This&That (Wang et al. 2025a). The DiT version further elevates performance, attaining 93.98 FVD and 0.847 SSIM with notably precise trajectory adherence (15.38 robot, 16.59 object error).
@@
VBench evaluation (Tab. 3) reveals our method’s balanced strengths: the SVD model leads in temporal stability (97.98 flickering score) and consistency metrics (95.39 subject, 96.57 background), while the DiT variant maintains competitive Imaging quality (68.65) with robust motion handling (98.70 smoothness). These results collectively validate our approach’s effectiveness in both visual quality and trajectory fidelity. We also visualize some of the same examples generated by our ManipDreamer3D based on SVD, and This&That are shown in Fig. 5. As shown in the figure, our model keeps the original shape of the original objects, while This&That suffers from object deformation. Fig. 5 presents a side-by-side comparison of manipulation sequences generated by our SVD-based ManipDreamer3D and This&That. The visual results demonstrate our method’s superior ability to maintain object structural integrity throughout the manipulation process, while This&That suffers from progressive object deformation that deviates from the original geometry. This qualitative advantage aligns with our quantitative metrics, showing better visual quality.
Figure 5: Qualitative comparison between our ManipDreamer3D and This&That (both SVD-based). Our method better preserves object appearance compared to This&That, which exhibits noticeable shape distortions in manipulation results.
Figure 6: Video visualization of path optimization effect.
Table 4: The video metric result of videos generated using the raw path and the optimized path.
Method Aesthetic Imaging Temporal Motion Subject background
w/o trajectory optimization 53.29 66.03 96.93 97.81 95.13 96.07
w/ trajectory optimization 53.67 66.51 96.89 97.81 95.16 96.08
@@
To evaluate the impact of trajectory optimization, we compare videos generated using initial path $P^{3}_{init}$ and optimized path $P^{3}_{opt}$ from our path planning method. As shown in Fig. 6, the $P^{3}_{opt}$ trajectory safely avoids the sink edge by moving upward, while $P^{3}_{init}$ produces an unsafe path along the sink edge. Quantitative results in Tab. 4 demonstrate that $P^{3}_{init}$ leads to worse video quality. The direct usage of videos generated with initial trajectory may cause unsafe behaviors in downstream VLA applications, confirming the importance of trajectory optimization.
Precision Manipulation Control
To evaluate the model’s capability in precision manipulation control, we conduct experiments using different functional parts (affordances) of objects to generate manipulation videos. The experiments employ our SVD-based model architecture to demonstrate fine-grained control over manipulation actions.
As illustrated in Fig. 7, we systematically vary the contact points on a pot object to test the model’s responsiveness to different manipulation conditions. The results show that our model successfully generates corresponding manipulation videos that accurately follow the specified affordance conditions, demonstrating its ability to handle intricate manipulation tasks.
@@ Conclusion
We propose ManipDreamer3D , a framework for generating robotic manipulation videos guided by 3D occupancy-aware trajectories. First, we reconstruct the scene in 3D, then plan efficient end-effector and object trajectories, and finally synthesize a coherent video from the first-frame latent. This design enables comprehensive control at keypoint, full-trajectory, and affordance levels while requiring minimal manual annotation. Extensive experiments on diverse scenes demonstrate that ManipDreamer3D  improves visual quality and spatial consistency, achieving stronger trajectory adherence compared with prior motion-controlled video generators.
@@
Our current planning primarily targets rigid-body interactions and quasi-static grasps. Future work will explore more contact- and compliance-aware objectives, together with trajectory-conditioned generative priors that better capture articulation and deformation.

### v2

@@ Introduction
To evaluate the quality of generated videos, we use common video qualities including FVD (unterthiner2018towards), SSIM (wang2004image), and PSNR (hore2010image), as well as VBench (huang2024vbench) metrics. Aside from that, we also evaluate our model with trajectory error, which reveals how well the model follows the trajectory. Following  (fu2025learning), we evaluated this metric for the robot end-effector and the target object, respectively.
Figure 5: Qualitative comparison between our ManipDreamer3D and This&That (both SVD-based). Our method better preserves object appearance compared to baseline This&That, which exhibits noticeable shape distortions in manipulation results.
Figure 6: Video synthesis comparison between using initial trajectories and optimized trajectories.
@@
The quantitative comparisons of video generation quality are presented in Tab. 2 and Tab. 3. Our ManipDreamer3D demonstrates superior performance across both evaluation frameworks. In common video metrics (Tab. 2), our SVD variant achieves best-in-class scores with FVD (143.33), PSNR (22.75), and SSIM (0.807), while significantly reducing trajectory errors (17.40 for robot, 18.77 for object) compared to SVD-based competitors including DragAnything (wu2024draganything) and This&That (wang2024language). The DiT version further elevates performance, attaining 93.98 FVD and 0.847 SSIM with notably precise trajectory adherence (15.38 for robot and 16.59 for object).
@@
VBench evaluation results in Tab. 3 reveals our method’s balanced strengths: the SVD model leads in temporal stability (97.98 flickering score) and consistency metrics (95.39 subject, 96.57 background), while the DiT variant maintains competitive Imaging quality (68.65) with robust motion handling (98.70 smoothness). These results collectively validate our approach’s effectiveness in both visual quality and trajectory fidelity. We also visualize some of the same examples generated by our ManipDreamer3D (SVD) and This&That. As shown in Fig. 5, our model keeps the original shape of the original objects, while This&That suffers from object deformation. This qualitative advantage aligns with our quantitative metrics, showing our superior visual quality.
@@
Precision Manipulation Control
To evaluate the model’s capability in precision manipulation control, we conduct experiments using different grasp parts (affordances) of objects to generate manipulation videos. The experiments employ our SVD-based model architecture to demonstrate fine-grained control over manipulation actions.
As illustrated in Fig. 7, we systematically vary the contact points on a pot object to test the model’s responsiveness to different manipulation positions. The results show that our model successfully generates manipulation videos that accurately follow the specified conditions, demonstrating its ability to handle intricate manipulation tasks.
@@
To evaluate the impact of trajectory optimization, we compare videos generated using initial path $P^{3}_{init}$ and optimized path $P^{3}_{opt}$ by our optimization method. As shown in Fig. 6, with the $P^{3}_{opt}$ trajectory, the generated video safely avoids colliding by moving upward, while $P^{3}_{init}$ produces an possibly unsafe path close to the sink edge. Direct usage of videos from initial trajectory may cause unsafe behaviors in downstream VLA models, confirming the importance to optimize trajectories.
@@ Conclusion
We propose ManipDreamer3D , a framework for generating robotic manipulation videos guided by 3D occupancy-aware trajectories. First, we reconstruct the scene in 3D, then plan efficient gripper and object trajectories, and finally synthesize a coherent video from the first-frame latent. This design enables comprehensive control at keypoint, full-trajectory, and affordance levels while requiring minimal manual annotation. Extensive experiments on diverse scenes demonstrate that ManipDreamer3D improves visual quality and spatial consistency, achieving stronger trajectory adherence compared with prior motion-controlled video generators.
@@
Our current planning primarily targets rigid-body interactions and quasi-static grasps. Future work will explore more contact- and compliance-aware objectives, together with generative priors that better capture articulation and deformation.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62476011).